A 5000 KG TRUCK IS RUNNING AT 72 KPH.


a. ) IF IT IS TO BE STOPPED BY A CONSTANT RETARDING


FORCE IN 10 S. WHAT IS THE DECELERATION ?


b ) WHAT IS THE RETARDING FORCE ?

1
Expert's answer
2022-02-15T09:58:27-0500

Given:

"m=\rm 5000\: kg"

"v_i=\rm 20\: m/s"

"v_f=\rm 0\: m/s"

"\Delta t=\rm 10\: s"


(a) the deceleration of the truck

"a=\frac{\Delta v}{\Delta t}=\frac{-20}{10}=\rm -2\: m/s^2"

(b) the retarding force

"F=ma=5000*2=10000\:\rm N"